Suppose the lengths of the pregnancies of a certain animal are approximately normally distributed with mean mu equals 281 days μ=281 days and standard deviation sigma equals 28 days.

(b) Suppose a random sample of 21 pregnancies is obtained. Describe the sampling distribution of the sample mean length of pregnancies.

The sampling distribution of x is _____ (skewed left, skewed right, normal) with u-x = _____ and o-x= _____. (Round to 4 decimal places as needed)
Answer #1

Normal distribution with,

  \mu = 281

\sigma = 28

A sample of size n = 21 is taken from this population.

Here , although the sample size is small , the sampling distribution of the sample mean is normal because the population distribution is normal.

The sampling distribution of the \bar x is normal with

\mu_{\bar x} = \mu = 281 and \sigma_{\bar x} =    \frac{\sigma}{\sqrt{n}} = 28/\sqrt{}​21 = 6.1101

Answer :

The sampling distribution of the \bar x is normal with  \mu_{\bar x} =  281 and \sigma_{\bar x} = 6.1101